OpenStudy (anonymous):

Solve the equation using the addition method (fraction) 15/7 x + 5y = -5 3x + 7y = -7

OpenStudy (solomonzelman):

\(\large\color{slate}{ \frac{\LARGE 15}{\LARGE 7} x + 5y = -5 }\) \(\large\color{slate}{ 3x + 7y = -7 }\) 1) Divide the first equation by 5. 2) multiply the first equation times -7. 3) add the equations. (this is going to eliminate the y for you

OpenStudy (anonymous):

so what is 15/7 divided by 5?

OpenStudy (anonymous):

i just get a long decimal

OpenStudy (anonymous):

@SolomonZelman ????

OpenStudy (solomonzelman):

15 / 5 =?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so 3/7?

OpenStudy (solomonzelman):

yes, if you divide 15/7 by 5 you get 3/7

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

so your new first equation, after dividing it by 5 will be?

OpenStudy (anonymous):

so it would be 3/7 + y = -1 divide that by -7 and it's....

OpenStudy (anonymous):

? + -7y = 7

OpenStudy (anonymous):

3/7 divided by -7 is____?

OpenStudy (anonymous):

@SolomonZelman ?

OpenStudy (solomonzelman):

you get after dividing by 5, 3/7 x + y = -1 now mltiply times -7

OpenStudy (anonymous):

3 + -7y = 7

OpenStudy (anonymous):

now what?

OpenStudy (solomonzelman):

3/7 * -7 = -3 1 * -7 = -7 -1 * -7 = 7 So, 3/7 x * -7 = -3x y * -7 = -7y -1 * -7 = 7 So your 1st equation would now be: \(\large\color{teal}{ -3x - 7y = 7 }\) \(\large\color{teal}{ 3x + 7y = -7 }\)

OpenStudy (anonymous):

Ok, what do I do once I have the first equation?

OpenStudy (solomonzelman):

if you have multiplied the first equation times 7, instead of multiplying times -7, then it would be: \(\large\color{teal}{ 3x + 7y = -7 }\) and your second equation as it was, is, \(\large\color{teal}{ 3x + 7y = -7 }\)

OpenStudy (solomonzelman):

both equations are same, no?

OpenStudy (anonymous):

so the answer is just 3x + 7y = -7? I need two numbers in a ( , ) thingy

OpenStudy (solomonzelman):

when two equations are same, the have infinite solutions.

OpenStudy (anonymous):

Ohhh ok. Thank you so much for the help! :)

OpenStudy (solomonzelman):

sure